AFTER SOLVING A SYSTEM OF LINEAR EQUATIONS, HOW CAN YOU CHECK YOUR ANSWER TO DETERMINE WHETHER IT IS CORRECT? please demonstrate the process by checking an answer for a set of equations (that you have created)
THANK YOU !

Answer by mananth(16946)

3x+2y=4 .............1
5x+2y=8 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-3x-2y= -4
5x+2y=8
Add the two equations
2x=4
/2
x=2
plug value of x in (1)
3x+2y=4
6+2y=4
2y=4-6
2y=-2
y=-1
CHECK
plug the value of x & y in any equation from above

3x+2y=4
6-2=4
4=4